How do you solve 1/5(a+10)=-3?
Answer by gabbymad98(6)

1/5 (a + 10) = -3 (Mutiply 1/5 to a and 10)
1/5a + 2 = -3 (Subtract 2 from -3)
1/5a = -5 (Divide 1/5 by -5, 1/5 divided by 1/5 should cancel out)
a = -25
Check:
1/5 (-25 + 10) = -3 (Multiply 1/5 to -25 and 10)
-5 + 2 = -3 (Add 2 to -5)
-3 = -3
Answer: a = -25